Consecutive problem:
If three times the smaller of two consecutive even integers is added to the larger, the result is 4 more than three times the larger. Find the two numbers (solve algebraically).
This is what I have:
3x+x=4x+3x

Found 2 solutions by mananth, amoresroy:
Answer by mananth(16949)

let the integers be x & x+2
..
3x+x+2 =3(x+2)+4
4x+2=3x+6+4
4x-3x=6+4-2
x=8
the integers are
8, 10
...
m.ananth@hotmail.ca

Answer by amoresroy(361)

If three times the smaller of two consecutive even integers is added to the larger, the result is 4 more than three times the larger. Find the two numbers (solve algebraically).
This is what I have:
3x+x=4x+3x
Let x = the smaller even integer
x+2= the larger even integer
3x + (x+2) = 3(x+2) + 4
4x + 2 = 3x + 6 + 4
x = 10-2
------------------------------
x = 8
x+2 = 10
------------------------------
The 2 numbers are 8 & 10